ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 261, 2024

Ancona, M. and A. Lerario Symplectic instability of Bézout’s theorem

841

Ash, D. D. and N. S. Ormes Topological speedups for minimal Cantor systems

91

Basok, M., S. O. Ivanov and R. Mikhailov Homology of the pronilpotent completion and cotorsion groups

281

Bering IV, E. A. and D. Studenmund Hall’s universal group is a subgroup of the abstract commensurator of a free group

493

di Blasio, G., F. Feo and G. Zecca Regularity results for local solutions to some anisotropic elliptic equations

1

Buser, P., E. Makover and B. Muetzel Short homology bases for hyperelliptic hyperbolic surfaces

977

Carrasco, P. D. and F. Rodriguez-Hertz Contributions to the ergodic theory of hyperbolic flows: unique ergodicity for quasiinvariant measures and equilibrium states for the time-one map

589

Casazza, P. G., S. J. Dilworth, D. Kutzarova and P. Motakis On uniqueness and plentitude of subsymmetric sequences

613

Cheng, Y. On the Rankin–Selberg L-factors for SO_{5} × GL_{2}  313

Davis, R., R. Pries and K. Wickelgren The Galois action on the lower central series of the fundamental group of the Fermat curve

171

Dilworth, S. J. See P. G. Casazza Feo, F. See G. di Blasio Fletcher, A. N. and J. Pratscher On the mean radius of quasi-

conformal mappings

419

Fonarev, A. and D. Kaledin Bökstedt periodicity generator via

K-theory 695

Garonzi, M., Z. Halasi and G. Somlai On the diameter of Cayley graphs of classical groups with generating sets containing a transvection

899

Gonçalves, D. L. and J. Guaschi Orbit configuration spaces and the homotopy groups of the pair (∏_{1}^{n} M, F_{n}(M)) for M either 𝕊^{2}

or ℝP^{2} 37

Greenberg, N. and S. Shelah Many forcing axioms for all regular uncountable cardinals

127

Guaschi, J. See D. L. Gonçalves Halasi, Z. See M. Garonzi

Hirose, M. Modular phenomena for regularized double zeta values  501
Humphries, P. The newform K-type and p-adic spherical harmonics  357

Ivanov, S. O. See M. Basok

Joó, A. On the packing/covering conjecture of infinite matroids  765

Kaledin, D. See A. Fonarev Kojman, M., A. Rinot and J. Steprāns Ramsey theory over par-

titions I: Positive Ramsey relations from forcing axioms

205

Kojman, M., A. Rinot and J. Steprāns Ramsey theory over partitions II: Negative Ramsey relations and pump-up theorems

223

Kutzarova, D. See P. G. Casazza Lerario, A. See M. Ancona

Li, H. and K. Liu Bernoullicity of lopsided principal algebraic actions  739

Liu, K. See H. Li Makover, E. See P. Buser Mikhailov, R. See M. Basok Mildenberger, M. L. a. H. Distributivity and minimality in perfect

tree forcings for singular cardinals

549

Möller, P., L. Paris and O. Varghese On parabolic subgroups of Artin groups

809

Mono, A. Locally harmonic Maaß forms of positive even weight  671

Motakis, P. See P. G. Casazza Muetzel, B. See P. Buser

Nenadov, R. Probabilistic hypergraph containers  879

de Oliveira, M. A., R. B. dos Santos and A. C. Vieira Polynomial growth of the codimensions sequence of algebras with group graded involution

445

Ormes, N. S. See D. D. Ash Paredes, M. and R. Sasyk Effective Hilbert’s irreducibility theorem

for global fields

851

Paris, L. See P. Möller Pratscher, J. See A. N. Fletcher Pries, R. See R. Davis

Ramaré, O. Notes on restriction theory in the primes  717

Rinot, A. See M. Kojman Rodriguez-Hertz, F. See P. D. Carrasco dos Santos, R. B. See M. A. de Oliveira Sasyk, R. See M. Paredes Shaul, L. and J. Williamson Lifting (co)stratifications between

tensor triangulated categories

249

Shelah, S. See N. Greenberg Slomka, B. A. A remark on discrete Brunn–Minkowski type inequal-

ities via transportation of measure

791

Somlai, G. See M. Garonzi

Stelzer, M. The cube axiom and resolutions in homotopy theory  395

Steprāns, J. See M. Kojman Studenmund, D. See E. A. Bering IV Tsui, M. C. Simplifying matrix differential equations with general co-

efficients

473

Varghese, O. See P. Möller Vieira, A. C. See M. A. de Oliveira

Wang, M. Waring–Goldbach problem in short intervals  637

Wickelgren, K. See R. Davis Williamson, J. See L. Shaul Yokoyama, T. Topological characterizations of recurrence, Poisson

stability, and isometric property of flows on surfaces

951

Zecca, G. See G. di Blasio